# On tori periods of Weil representations of unitary groups

### v1

@@ 3.2.1. Splitting set up
Here $\operatorname{U}(\mathrm{W}_{F,1}^{\delta})$ is realized as a subgroup of $\operatorname{U}(\mathrm{W}_{E,1}^{\delta})$ diagonally, that is, an element $x\in K^{1}\cong\operatorname{U}(\mathrm{W}_{F,1}^{\delta})$ acts on $\mathrm{W}_{E,1}^{\delta}$ by the multiplication by $x$ map $\mathrm{m}_{x}$, given by $\mathrm{m}_{x}w=xw$, for $w\in\mathrm{W}_{E,1}^{\delta}$. In order to write down a seesaw identity, we need to fix compatible splittings. We explain this now. Given characters $\chi_{L_{E,\lambda}},\chi_{\mathrm{W}_{E,1}^{\delta}}\colon L_{E}^{\times}\to\mathbb{C}^{\times}$, such that
@@ 3.2.2. The local seesaw identity
Let $\alpha\colon\operatorname{U}(L_{E,\lambda})\rightarrow\mathbb{C}^{\times}$ and $\beta\colon\operatorname{U}(\mathrm{W}_{F,1}^{\delta})\to\mathbb{C}^{\times}$ be characters. Choose compatible splittings $\tilde{\iota}$ and $\tilde{\iota}^{\prime}$ as above. Let $\Theta\left(\beta\right)$ be the big theta lift of $\beta$ to $\operatorname{U}(\mathrm{V}_{E,\lambda})$ with respect to the splitting $\tilde{\iota}^{\prime}$. By (4), we have the following equality, where we take $H=\operatorname{U}(L_{E,\lambda})$:
@@ 3.3.1. Vector of root numbers of characters of $L_{E}^{\times}$
Let $\psi\colon F\rightarrow\mathbb{C}^{\times}$ be a non-trivial character. For any trace zero element $\delta\in K^{\times}$ we define $\psi_{\delta}\colon K\rightarrow\mathbb{C}^{\times}$ by the formula $\psi_{\delta}\left(x\right)=\psi\left(\operatorname{tr}_{K/F}(\delta x)\right)$. Note that for any $x\in K^{\times}$, we have $\psi_{\delta}\left(x^{c}\right)=\psi_{\delta}^{-1}\left(x\right)$. For any character $\chi\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ such that $\chi\left(x^{c}\right)=\chi\left(x^{-1}\right)$, we will define a vector of root numbers $\varepsilon\left(\chi,\psi,\delta\right)$. Assume first that $E$ is a field extension of $F$. The étale algebra $L_{E}=E\otimes_{F}K$ is either a field or is isomorphic to $E\times E$. Let $\chi\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ be a character as above. If $L_{E}$ is a field, we set
@@
where for a complex number $s$, the factor $\varepsilon^{\mathrm{Tate}}_{L_{E}}\left(s,\chi,\psi_{\delta}\circ\operatorname{tr}_{L_{E}/K}\right)$ is the epsilon factor defined by Tate [29, 20]. If $L_{E}=E\times E$, then we define $\varepsilon_{L_{E}/E}\left(\chi,\psi,\delta\right)=1$. We move to the general case. Let $E$ be a finite-dimensional étale algebra. As before, we may write $E=\prod_{j=1}^{m}F_{j}$, where $F_{j}/F$ is a field extension. Given a character $\chi\colon L_{E}^{\times}\to\mathbb{C}^{\times}$, we may regard it as a tuple $\left(\chi_{1},\dots,\chi_{m}\right)$, where $\chi_{j}\colon L_{F_{j}}^{\times}\rightarrow\mathbb{C}^{\times}$ is a character satisfying $\chi_{j}\left(x^{c}\right)=\chi_{j}^{-1}\left(x\right)$ for every $j$ and $x\in L_{F_{j}}^{\times}$. We define $\varepsilon_{L_{E}/E}\left(\chi,\psi,\delta\right)$ as the following tuple:
@@ 3.3.2. Base change for characters of $L_{E}^{1}$
We have an isomorphism $j_{E}\colon L_{E}^{\times}/E^{\times}\rightarrow L_{E}^{1}$ given by $j_{E}\left(x\right)=\frac{x}{x^{c}}$. Given a character $\beta\colon L_{E}^{1}\rightarrow\mathbb{C}^{\times}$, we define a character $\beta_{L_{E}}\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ by the formula
@@ 3.3.3. Non-vanishing of theta lifts
Suppose that $K/F$ is a quadratic field extension, and let $\delta\in K^{\times}$ be a trace zero element. Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate one-dimensional hermitian and skew hermitian spaces over $K$, respectively. The non-vanishing of a theta lift of a character of $\operatorname{U}(\mathrm{V})$ is treated separately for the archimedean case and the non-archimedean case. When $F$ is non-archimedean, the statement is given by [15, Theorem 6.1]. There are subtle differences between the versions of this result presented in [15] and the version we state below. Such differences are explained in [8, Section 9]. When $F=\mathbb{R}$, the result is given by [24, Theorem 6.1]. Once again, it is written in a different language, and we refer to [33, Section 3.2] for the translation [^1] We warn the reader that the characters $\chi_{\mathrm{V}}$ and $\chi_{\mathrm{V}^{\prime}}$ in [33] are $\chi_{\mathrm{W}}$ and $\chi_{\mathrm{V}}$, respectively, in our notation.. Let $i^{\prime}_{\mathrm{V}}\colon\operatorname{U}(\mathrm{V})\rightarrow K^{1}$ and $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\rightarrow K^{1}$ be admissible isomorphisms. Let $\alpha\colon K^{1}\to\mathbb{C}^{\times}$ be a character. The following result determines the theta lift $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\alpha\circ i^{\prime}_{\mathrm{V}}\right)$.
@@
The theta lift $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\alpha\circ i^{\prime}_{\mathrm{V}}\right)$ with respect to the splitting $\tilde{\iota}$ associated to characters $(\chi_{\mathrm{V}},\chi_{\mathrm{W}})$ is non-zero if and only if
@@ 3.3.4. Non-vanishing of theta lifts for one-dimensional spaces over an étale algebra
Let $E$ be an étale algebra of degree $n$ over $F$. Let $\lambda\in E^{\times}$ and let $\delta\in K^{\times}$ be a trace zero element. Consider the same setup as in Section 3.2. Theorem 3.4 has an obvious extension for the spaces $\mathrm{W}_{E,1}^{\delta}$ and $\mathrm{V}_{E,\lambda}$, which will be useful for the proof of our main result. Let $i^{\prime}_{\mathrm{V}}\colon\operatorname{U}(L_{E,\lambda})\rightarrow L_{E}^{1}$ and $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W}_{E,1}^{\delta})\rightarrow L_{E}^{1}$ be admissible isomorphisms. Let $\alpha\colon L_{E}^{1}\rightarrow\mathbb{C}^{\times}$ be a character.
@@ 4.1. The global theta correspondence
Let $\mathbf{F}$ be a number field and let $\mathbf{K}/\mathbf{F}$ be a quadratic field extension with involution $x\mapsto x^{c}$, whose set of fixed points is $\mathbf{F}$. Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate finite dimensional hermitian and skew-hermitian spaces over $\mathbf{K}$, respectively. As in the local case, we consider the tensor product $\operatorname{Res}_{\mathbf{K}/\mathbf{F}}\left(\mathrm{V}\otimes_{\mathbf{F}}\mathrm{W}\right)$. Let $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)=\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}/\mathbf{F}}\left(\mathrm{V}\otimes_{\mathbf{K}}\mathrm{W}\right)\right)$. Let $\mathbb{A}_{\mathbf{F}}$ be the adeles of $\mathbf{F}$. For an algebraic group $G$ defined over $\mathbf{F}$, denote by $\left[G\right]=G\left(\mathbf{F}\right)\backslash G\left(\mathbb{A}_{\mathbf{F}}\right)$ its automorphic quotient. Let $\psi:\mathbf{F}\backslash\mathbb{A}_{\mathbf{F}}\rightarrow\mathbb{C}^{\times}$ be a non-trivial character. Write $\psi=\bigotimes_{v}\psi_{v}$. For every place $v$ of $\mathbf{F}$, we denote $\mathrm{V}_{v}=\mathrm{V}\otimes_{\mathbf{F}}\mathbf{F}_{v}$ and $\mathrm{W}_{v}=\mathrm{W}\otimes_{\mathbf{F}}\mathbf{F}_{v}$. We also denote $\mathbf{K}_{v}=\mathbf{K}\otimes_{\mathbf{F}}\mathbf{F}_{v}$. For almost all $v$, the covering
@@
splits uniquely over the maximal hyperspecial subgroup $\mathcal{K}_{v}$ of $\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}_{v}/\mathbf{F}_{v}}\left(\mathrm{V}_{v}\otimes_{\mathbf{K}_{v}}\mathrm{W}_{v}\right)\right)$. Let $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)=\prod_{v}^{\prime}\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}_{v}/\mathbf{F}_{v}}\left(\mathrm{V}_{v}\otimes_{\mathbf{K}_{v}}\mathrm{W}_{v}\right)\right)$ be the restricted product with respect to $\mathcal{K}_{v}\subset\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}_{v}/\mathbf{F}_{v}}\left(\mathrm{V}_{v}\otimes_{\mathbf{K}_{v}}\mathrm{W}_{v}\right)\right)$. Consider the restricted product $\prod^{\prime}_{v}\operatorname{Mp}_{\psi_{v}}\left(\mathrm{V}_{v},\mathrm{W}_{v}\right)$ with respect to $\mathcal{K}_{v}\subset\operatorname{Mp}_{\psi_{v}}\left(\mathrm{V}_{v},\mathrm{W}_{v}\right)$. We denote by $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ the quotient of the latter restricted product by the central subgroup
@@
We denote $\left[\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\right]=\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)\backslash\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$. Let $\operatorname{Res}_{\mathbf{K}/\mathbf{F}}\left(\mathrm{V}\otimes_{\mathbf{K}}\mathrm{W}\right)=\mathrm{X}\oplus\mathrm{Y}$ be a polarization, and for every $v$ let $\mathrm{X}_{v}=\mathrm{X}\otimes_{\mathbf{F}}\mathbf{F}_{v}$ and $\mathrm{Y}_{v}=\mathrm{Y}\otimes_{\mathbf{F}}\mathbf{F}_{v}$. For every $v$, we realize the Weil representation $\omega_{\psi_{v},\mathbf{F}_{v}}$ of $\operatorname{Mp}_{\psi_{v}}\left(\mathrm{V}_{v},\mathrm{W}_{v}\right)$ via its Schrödinger model, acting on the space $\mathcal{S}\left(\mathrm{Y}_{v}\right)$ consisting of Schwartz functions on $\mathrm{Y}_{v}$. We denote by $\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)=\bigotimes^{\prime}_{v}\mathcal{S}\left(\mathrm{Y}_{v}\right)$ the restricted tensor product and by $\omega_{\psi,\mathbb{A}_{\mathbf{F}}}=\bigotimes_{v}\omega_{\psi_{v},\mathbf{F}_{v}}$ the global Weil representation of $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$. For a function $\varphi\in\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)$, we consider its theta series, defined for $g\in\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ by
@@
Then it is well-known that $\theta\left(\varphi\right)$ is an automorphic form of $\left[\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\right]$. The embedding $\iota\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\to\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)$ discussed in Section 3.1 has a global analog
@@
such that the image of $\operatorname{U}(\mathrm{V})\left(\mathbf{F}\right)\times\operatorname{U}(\mathrm{W})\left(\mathbf{F}\right)$ under $\tilde{\iota}$ lies in $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)$. Such a lifting exists, and we postpone the discussion regarding the data needed in order to construct it to the next subsection. For a cuspidal automorphic form $f:\left[\operatorname{U}(\mathrm{V})\right]\to\mathbb{C}$, a Schwartz function $\varphi\in\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)$, and an element $g_{\mathrm{W}}\in\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right),$ we denote
@@
and call $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ the global theta lift of $\pi$. By [21, Corollary 7.3] if $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ lies in the space of square-integrable automorphic forms, then $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)=\bigotimes^{\prime}_{v}\theta_{\mathrm{V}_{v},\mathrm{W}_{v},\tilde{\iota}_{v},\psi_{v}}\left(\pi_{v}\right)$.

### v2

@@ 3.2.1. Splitting set up
Here $\operatorname{U}(\mathrm{W}_{F,1}^{\delta})$ is realized as a subgroup of $\operatorname{U}(\mathrm{W}_{E,1}^{\delta})$ diagonally, that is, an element $x\in K^{1}\cong\operatorname{U}(\mathrm{W}_{F,1}^{\delta})$ acts on $\mathrm{W}_{E,1}^{\delta}$ by the multiplication by $x$ map $\mathrm{m}_{x}$, given by $\mathrm{m}_{x}w=xw$, for $w\in\mathrm{W}_{E,1}^{\delta}$.
In order to write down a seesaw identity, we need to fix compatible splittings. We explain this now.
Given characters $\chi_{L_{E,\lambda}},\chi_{\mathrm{W}_{E,1}^{\delta}}\colon L_{E}^{\times}\to\mathbb{C}^{\times}$, such that
@@ 3.2.2. The local seesaw identity
Let $\alpha\colon\operatorname{U}(L_{E,\lambda})\rightarrow\mathbb{C}^{\times}$ and $\beta\colon\operatorname{U}(\mathrm{W}_{F,1}^{\delta})\to\mathbb{C}^{\times}$ be characters.
Choose compatible splittings $\tilde{\iota}$ and $\tilde{\iota}^{\prime}$ as above. Let $\Theta\left(\beta\right)$ be the big theta lift of $\beta$ to $\operatorname{U}(\mathrm{V}_{E,\lambda})$ with respect to the splitting $\tilde{\iota}^{\prime}$. By (4), we have the following equality, where we take $H=\operatorname{U}(L_{E,\lambda})$:
@@ 3.3.1. Vector of root numbers of characters of $L_{E}^{\times}$
Let $\psi\colon F\rightarrow\mathbb{C}^{\times}$ be a non-trivial character. For any trace zero element $\delta\in K^{\times}$ we define $\psi_{\delta}\colon K\rightarrow\mathbb{C}^{\times}$ by the formula $\psi_{\delta}\left(x\right)=\psi\left(\operatorname{tr}_{K/F}(\delta x)\right)$. Note that for any $x\in K^{\times}$, we have $\psi_{\delta}\left(x^{c}\right)=\psi_{\delta}^{-1}\left(x\right)$.
For any character $\chi\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ such that $\chi\left(x^{c}\right)=\chi\left(x^{-1}\right)$, we will define a vector of root numbers $\varepsilon\left(\chi,\psi,\delta\right)$.
Assume first that $E$ is a field extension of $F$. The étale algebra $L_{E}=E\otimes_{F}K$ is either a field or is isomorphic to $E\times E$. Let $\chi\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ be a character as above. If $L_{E}$ is a field, we set
@@
where for a complex number $s$, the factor $\varepsilon^{\mathrm{Tate}}_{L_{E}}\left(s,\chi,\psi_{\delta}\circ\operatorname{tr}_{L_{E}/K}\right)$ is the epsilon factor defined by Tate [30, 20]. If $L_{E}=E\times E$, then we define $\varepsilon_{L_{E}/E}\left(\chi,\psi,\delta\right)=1$.
We move to the general case. Let $E$ be a finite-dimensional étale algebra. As before, we may write $E=\prod_{j=1}^{m}F_{j}$, where $F_{j}/F$ is a field extension. Given a character $\chi\colon L_{E}^{\times}\to\mathbb{C}^{\times}$, we may regard it as a tuple $\left(\chi_{1},\dots,\chi_{m}\right)$, where $\chi_{j}\colon L_{F_{j}}^{\times}\rightarrow\mathbb{C}^{\times}$ is a character satisfying $\chi_{j}\left(x^{c}\right)=\chi_{j}^{-1}\left(x\right)$ for every $j$ and $x\in L_{F_{j}}^{\times}$. We define $\varepsilon_{L_{E}/E}\left(\chi,\psi,\delta\right)$ as the following tuple:
@@ 3.3.2. Base change for characters of $L_{E}^{1}$
We have an isomorphism $j_{E}\colon L_{E}^{\times}/E^{\times}\to L_{E}^{1}$ given by $j_{E}\left(x\right)=\frac{x}{x^{c}}$.
Given a character $\beta\colon L_{E}^{1}\rightarrow\mathbb{C}^{\times}$, we define a character $\beta_{L_{E}}\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ by the formula
@@ 3.3.3. Non-vanishing of theta lifts
Suppose that $K/F$ is a quadratic field extension, and let $\delta\in K^{\times}$ be a trace zero element. Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate one-dimensional hermitian and skew hermitian spaces over $K$, respectively.
The non-vanishing of a theta lift of a character of $\operatorname{U}(\mathrm{V})$ is treated separately for the archimedean case and the non-archimedean case. When $F$ is non-archimedean, the statement is given by [15, Theorem 6.1]. There are subtle differences between the versions of this result presented in [15] and the version we state below. Such differences are explained in [8, Section 9]. When $F=\mathbb{R}$, the result is given by [25, Theorem 6.1]. Once again, it is written in a different language, and we refer to [34, Section 3.2] for the translation [^1] We warn the reader that the characters $\chi_{\mathrm{V}}$ and $\chi_{\mathrm{V}^{\prime}}$ in [34] are $\chi_{\mathrm{W}}$ and $\chi_{\mathrm{V}}$, respectively, in our notation.. Let $i^{\prime}_{\mathrm{V}}\colon\operatorname{U}(\mathrm{V})\to K^{1}$ and $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to K^{1}$ be the obvious isomorphisms. Let $\alpha\colon K^{1}\to\mathbb{C}^{\times}$ be a character. The following result determines the theta lift $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\alpha\circ i^{\prime}_{\mathrm{V}}\right)$.
@@
The theta lift $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\alpha\circ i^{\prime}_{\mathrm{V}}\right)$ with respect to the splitting $\tilde{\iota}$ associated to the characters $(\chi_{\mathrm{V}},\chi_{\mathrm{W}})$ is non-zero if and only if
@@ 3.3.4. Non-vanishing of theta lifts for one-dimensional spaces over an étale algebra
Let $E$ be an étale algebra of degree $n$ over $F$. Let $\lambda\in E^{\times}$ and let $\delta\in K^{\times}$ be a trace zero element. Consider the same setup as in Section 3.2. Theorem 3.5 has an obvious extension for the spaces $\mathrm{W}_{E,1}^{\delta}$ and $\mathrm{V}_{E,\lambda}$, which will be useful for the proof of our main result.
Let $i^{\prime}_{\mathrm{V}}\colon\operatorname{U}(L_{E,\lambda})\to L_{E}^{1}$ and $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W}_{E,1}^{\delta})\to L_{E}^{1}$ be the obvious isomorphisms. Let $\alpha\colon L_{E}^{1}\rightarrow\mathbb{C}^{\times}$ be a character.
@@ 4.1. The global theta correspondence
Let $\mathbf{F}$ be a number field and let $\mathbf{K}/\mathbf{F}$ be a quadratic field extension with involution $x\mapsto x^{c}$, whose set of fixed points is $\mathbf{F}$.
Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate finite dimensional hermitian and skew-hermitian spaces over $\mathbf{K}$, respectively. As in the local case, we consider the tensor product $\operatorname{Res}_{\mathbf{K}/\mathbf{F}}\left(\mathrm{V}\otimes_{\mathbf{F}}\mathrm{W}\right)$. Let $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)=\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}/\mathbf{F}}\left(\mathrm{V}\otimes_{\mathbf{K}}\mathrm{W}\right)\right)$.
Let $\mathbb{A}_{\mathbf{F}}$ be the adeles of $\mathbf{F}$. For an algebraic group $G$ defined over $\mathbf{F}$, denote by $\left[G\right]=G\left(\mathbf{F}\right)\backslash G\left(\mathbb{A}_{\mathbf{F}}\right)$ its automorphic quotient. Let $\psi:\mathbf{F}\backslash\mathbb{A}_{\mathbf{F}}\rightarrow\mathbb{C}^{\times}$ be a non-trivial character. Write $\psi=\bigotimes_{v}\psi_{v}$. For every place $v$ of $\mathbf{F}$, we denote $\mathrm{V}_{v}=\mathrm{V}\otimes_{\mathbf{F}}\mathbf{F}_{v}$ and $\mathrm{W}_{v}=\mathrm{W}\otimes_{\mathbf{F}}\mathbf{F}_{v}$. We also denote $\mathbf{K}_{v}=\mathbf{K}\otimes_{\mathbf{F}}\mathbf{F}_{v}$.
For almost all $v$, the covering
@@
splits uniquely over the maximal hyperspecial subgroup $\mathcal{K}_{v}$ of $\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}_{v}/\mathbf{F}_{v}}\left(\mathrm{V}_{v}\otimes_{\mathbf{K}_{v}}\mathrm{W}_{v}\right)\right)$. Let
$\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)={\prod_{v}}^{\prime}\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}_{v}/\mathbf{F}_{v}}\left(\mathrm{V}_{v}\otimes_{\mathbf{K}_{v}}\mathrm{W}_{v}\right)\right)$
be the restricted product with respect to $\mathcal{K}_{v}\subset\operatorname{Sp}\left(\operatorname{Res}_{\mathbf{K}_{v}/\mathbf{F}_{v}}\left(\mathrm{V}_{v}\otimes_{\mathbf{K}_{v}}\mathrm{W}_{v}\right)\right)$. Consider the restricted product $\prod^{\prime}_{v}\operatorname{Mp}_{\psi_{v}}\left(\mathrm{V}_{v},\mathrm{W}_{v}\right)$ with respect to $\mathcal{K}_{v}\subset\operatorname{Mp}_{\psi_{v}}\left(\mathrm{V}_{v},\mathrm{W}_{v}\right)$. We denote by $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ the quotient of the latter restricted product by the central subgroup
@@
We have that $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ splits canonically over $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)$. Thus, we may regard $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)$ as a subgroup of $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ and define
$\left[\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\right]=\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)\backslash\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right).$
Let $\operatorname{Res}_{\mathbf{K}/\mathbf{F}}\left(\mathrm{V}\otimes_{\mathbf{K}}\mathrm{W}\right)=\mathrm{X}\oplus\mathrm{Y}$ be a polarization, and for every $v$ let $\mathrm{X}_{v}=\mathrm{X}\otimes_{\mathbf{F}}\mathbf{F}_{v}$ and $\mathrm{Y}_{v}=\mathrm{Y}\otimes_{\mathbf{F}}\mathbf{F}_{v}$. For every $v$, we realize the Weil representation $\omega_{\psi_{v},\mathbf{F}_{v}}$ of $\operatorname{Mp}_{\psi_{v}}\left(\mathrm{V}_{v},\mathrm{W}_{v}\right)$ via its Schrödinger model, acting on the space $\mathcal{S}\left(\mathrm{Y}_{v}\right)$ consisting of Schwartz functions on $\mathrm{Y}_{v}$. We denote by $\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)=\bigotimes^{\prime}_{v}\mathcal{S}\left(\mathrm{Y}_{v}\right)$ the restricted tensor product and by $\omega_{\psi,\mathbb{A}_{\mathbf{F}}}=\bigotimes_{v}\omega_{\psi_{v},\mathbf{F}_{v}}$ the global Weil representation of $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$.
For a function $\varphi\in\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)$, we consider its theta series, defined for $g\in\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ by
@@
Then it is well-known that $\theta\left(\varphi\right)$ is an automorphic form of $\left[\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\right]$.
The embedding $\iota\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\to\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)$ discussed in Section 3.1 has a global analog
@@
such that the image of $\operatorname{U}(\mathrm{V})\left(\mathbf{F}\right)\times\operatorname{U}(\mathrm{W})\left(\mathbf{F}\right)$ under $\tilde{\iota}$ lies in $\operatorname{Sp}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbf{F}\right)$. Such a lifting exists, and we postpone the discussion regarding the data needed in order to construct it to the next subsection.
For a cuspidal automorphic form $f:\left[\operatorname{U}(\mathrm{V})\right]\to\mathbb{C}$, a Schwartz function $\varphi\in\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)$, and an element $g_{\mathrm{W}}\in\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right),$ we denote
@@
and call $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ the global theta lift of $\pi$.
By [21, Corollary 7.3] if $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ lies in the space of square-integrable automorphic forms, then $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)=\bigotimes^{\prime}_{v}\theta_{\mathrm{V}_{v},\mathrm{W}_{v},\tilde{\iota}_{v},\psi_{v}}\left(\pi_{v}\right)$.